# A Population Analysis of 20 Exoplanets Observed from the Optical of the Near-infrared Wavelengths with HST: Evidence for Widespread Stellar Contamination

### v1

@@ 1 Introduction
The paper is organised as follows: in Section 2 we describe the procedures we employed in our HST-dedicated data analysis pipeline to reduce and calibrate STIS and WFC3 data (Section 2.1 & 2.2), the light-curve detrending techniques applied to derive the final spectrum (Section 2.3) and a section on parameter estimation (Section 2.4) for completeness. The modelling methodology, including the variety of retrieval runs we performed on the data are described in Section 3. Section 4 describes our use of the Bayesian evidence/Bayes factor, a well-established statistical tool (Section 4.1) alongside two newly-developed metrics (Section 4.2, 4.3) that we employed to describe the level of stellar contamination across our planetary sample. The results of our investigation are reported in Section 5, while population level findings together with three distinct case-studies are explored in more detail in Section 6. Finally, a summary highlighting our discoveries and their implications is presented in Section 7.
@@ 2.1 STIS Data Processing
Starting from the raw STIS data, we performed data reduction and calibration following the procedures highlighted in the STIS documentation (Bostroem & Proffitt, 2011). The steps include: bias level subtraction (a.k.a. reference pixels correction), bias frame subtraction, dark image subtraction, background subtraction, flat-field correction, bad-pixel and cosmic-ray correction, wavelength calibration and light-curves extraction. For the extraction we used 6-pixel-wide apertures along the cross-dispersion direction and smoothed aperture edges along the dispersion direction (the smoothing factors were 5 Å for the G430L grating and 10 Å for the G750L grating, corresponding to approximately 2 pixels in each case). The spectral bin sizes were chosen in such a way that roughly the same amount of flux would fall in each of them while keeping the S/N approximately constant across the spectrum. For each planetary spectra we extracted a broad band (white) light-curve and a total of nine and seven spectral light-curves for G430L and G750L respectively. For a number of planets, some of the chromatic light-curves displayed strong outliers; we removed them by inspection before the light-curve fitting step.
@@ 2.2 WFC3 Data Processing
As highlighted in Tsiaras et al. (2016a), Iraclis starts from the raw data and performs a number of steps (zero-read subtraction, reference pixels correction, non-linearity correction, dark current subtraction, gain conversion, sky background subtraction, calibration, flat-field correction, bad-pixel and cosmic-ray correction) before extracting the light-curves. Both the white and spectral light-curves are then fitted with the PyLightcurve package embedded into the pipeline (more details in Section 2.3), to produce a final spectrum. Similarly to STIS, we chose the spectral bin sizes in such a way that roughly the same amount of flux would fall in each bin while keeping the S/N approximately constant across the whole spectrum.
2.3 Light-curve fitting
The light-curve modelling parameters were pulled from the Exoplanet Characterisation Catalogue developed as part of the ExoClock project (Kokori et al., 2021, 2022, 2023). In each of the light-curve modelling steps we assumed a circular orbit and a fixed period for the planet, while the limb-darkening effect was modelled using the Claret 4-coefficient law (Claret, 2000) and the ExoTETHyS (Morello et al., 2020) package, which takes into account the stellar parameters (Appendix C) and the response curve of the instrument. The white light-curves were fitted with a function M(t) that includes a transit model F(t) and a systematics model R(t) multiplied by a normalisation factor , of the form:
@@ 2.3 Light-curve fitting
While the white light-curves were fitted using the above technique for both WFC3 and STIS, we used two different fitting routines to treat WFC3 and STIS spectral light-curves. After excluding the initial orbit from each WFC3 spectral light-curve, as it exhibits more pronounced wavelength-dependent systematics compared to the subsequent orbits, we employed Equation 2 to model the remaining orbits comprising the spectral light-curves. The systematics-modelled light-curves were then divided by the corresponding white light-curve and further fitted via
where is a wavelength-dependent normalisation factor, is the coefficient of the wavelength-dependent linear slope, is time, is the mid-transit time, is the wavelength dependent transit model, and is the best-fit model on the white light-curve.
@@ 4.2 Stellar Activity Distance metric (SAD)
We select solely the data points 1.0 m since this is the wavelength range where stellar activity has the potential to induce the strongest chromatic effects. Due to the often contrasting STIS observations when multiple data sets are available for the same grating, we consider each Case on its own. For each Case we calculate the absolute difference of each observational data point from the stellar activity model and, separately, the absolute difference from the base model. We apply standard error propagation to calculate the uncertainty in the absolute difference, using the transit depth and model 1 errors.
The calculations effectively return an absolute difference and an associated error per data point related to both the base model and the stellar model. As an example, if we take the spectrum of WASP-127 b and we consider the 16 datapoints with wavelengths bluer than 1 m (see Fig. 2), we end up with sixteen absolute distances from the base model and sixteen absolute distances from the stellar model, for a total of thirty-two absolute distances. By considering each spectral Case separately, we then calculate the average of all distances from each model and its uncertainty, obtaining an average distance related to the base model () and one referring to the stellar model (). Finally, the stellar activity distance metric (SAD) is defined as the ratio between the two mean absolute distances
@@ Acknowledgements
Funding: This project has received funding from the European Research Council under the European Union Horizon 2020 research and innovation program (grant agreement No. 758892, ExoAI) and the European Union Horizon 2020 COMPET program (grant agreement No. 776403, ExoplANETS A). Furthermore, we acknowledge funding by the UK Space Agency and Science and Technology Funding Council grants ST/K502406/1, ST/P000282/1, ST/P002153/1, ST/ S002634/1, ST/T001836/1, ST/V003380/1, and ST/W00254X/1. This project has received funding from the European Research Council (ERC) under the
European Union Horizon 2020 research and innovation programme through Advanced Investigator grant number 883830 (ExoMolHD).
@@
Software: astropy (Astropy Collaboration et al., 2013, 2018; Price-Whelan et al., 2018),
Iraclis (Tsiaras et al., 2016b), PyLightcurve (Tsiaras et al., 2016a), ExoTETHyS (Morello et al., 2020), TauREx 3.1 (Al-Refaie et al., 2021), ASteRA (Thompson et al., 2024), YunMa (Ma et al., 2023), h5py (Collette, 2013), emcee (Foreman-Mackey et al., 2013), Matplotlib (Hunter, 2007), Multinest (Feroz et al., 2009), PyMultinest (Buchner et al., 2014), Numpy (Oliphant, 2006), corner (Foreman-Mackey, 2016).

### v2

# A Population Analysis of 20 Exoplanets Observed from the Optical to the Near-infrared Wavelengths with HST: Evidence for Widespread Stellar Contamination

@@ 1 Introduction
The paper is organised as follows: in Section 2 we describe the procedures we employed in our HST-dedicated data analysis pipeline to reduce and calibrate STIS and WFC3 data (Section 2.1 & 2.2), the light curve detrending techniques applied to derive the final spectrum (Section 2.3) and a section on parameter estimation (Section 2.4) for completeness. The modelling methodology, including the variety of retrieval runs we performed on the data are described in Section 3. Section 4 describes our use of the Bayesian evidence/Bayes factor, a well-established statistical tool (Section 4.1) alongside two newly-developed metrics (Section 4.2, 4.3) that we employed to describe the level of stellar contamination across our planetary sample. The results of our investigation are reported in Section 5, while population level findings together with three distinct case-studies are explored in more detail in Section 6. Finally, a summary highlighting our discoveries and their implications is presented in Section 7.
@@ 2.1 STIS Data Processing
Starting from the raw STIS data, we performed data reduction and calibration following the procedures highlighted in the STIS documentation (Bostroem & Proffitt, 2011). The steps include: bias level subtraction (a.k.a. reference pixels correction), bias frame subtraction, dark image subtraction, background subtraction, flat-field correction, bad-pixel and cosmic-ray correction, wavelength calibration and light curves extraction. For the extraction we used 6-pixel-wide apertures along the cross-dispersion direction and smoothed aperture edges along the dispersion direction (the smoothing factors were 5 Å for the G430L grating and 10 Å for the G750L grating, corresponding to approximately 2 pixels in each case). The spectral bin sizes were chosen in such a way that roughly the same amount of flux would fall in each of them while keeping the S/N approximately constant across the spectrum. For each planetary spectra we extracted a broad band (white) light curve and a total of nine and seven spectral light curves for G430L and G750L respectively. For a number of planets, some of the chromatic light curves displayed strong outliers; we removed them by inspection before the light curve fitting step.
@@ 2.2 WFC3 Data Processing
As highlighted in Tsiaras et al. (2016a), Iraclis starts from the raw data and performs a number of steps (zero-read subtraction, reference pixels correction, non-linearity correction, dark current subtraction, gain conversion, sky background subtraction, calibration, flat-field correction, bad-pixel and cosmic-ray correction) before extracting the light curves. Both the white and spectral light curves are then fitted with the PyLightcurve package embedded into the pipeline (more details in Section 2.3), to produce a final spectrum. Similarly to STIS, we chose the spectral bin sizes in such a way that roughly the same amount of flux would fall in each bin while keeping the S/N approximately constant across the whole spectrum.
2.3 Light curve fitting
The light curve modelling parameters were pulled from the Exoplanet Characterisation Catalogue developed as part of the ExoClock project (Kokori et al., 2021, 2022, 2023). In each of the light curve modelling steps we assumed a circular orbit and a fixed period for the planet, while the limb-darkening effect was modelled using the Claret 4-coefficient law (Claret, 2000) and the ExoTETHyS (Morello et al., 2020) package, which takes into account the stellar parameters (Appendix C) and the response curve of the instrument. The white light curves were fitted with a function M(t) that includes a transit model F(t) and a systematics model R(t) multiplied by a normalisation factor , of the form:
@@ 2.3 Light curve fitting
While the white light curves were fitted using the above technique for both WFC3 and STIS, we used two different fitting routines to treat WFC3 and STIS spectral light curves. After excluding the initial orbit from each WFC3 spectral light curve, as it exhibits more pronounced wavelength-dependent systematics compared to the subsequent orbits, we employed Equation 2 to model the remaining orbits comprising the spectral light curves. The systematics-modelled light curves were then divided by the corresponding white light curve and further fitted via
where is a wavelength-dependent normalisation factor, is the coefficient of the wavelength-dependent linear slope, is time, is the mid-transit time, is the wavelength dependent transit model, and is the best-fit model on the white light curve.
@@ 4.2 Stellar Activity Distance metric (SAD)
We select solely the data points 1.0 m since this is the wavelength range where stellar activity has the potential to induce the strongest chromatic effects. Due to the often contrasting STIS observations when multiple data sets are available for the same grating, we consider each Case on its own. For each Case we calculate the absolute difference of each observational data point from the stellar activity model and, separately, the absolute difference from the base model. We apply standard error propagation to calculate the uncertainty in the absolute difference, using the transit depth and model 1 errors. The calculations effectively return an absolute difference and an associated error per data point related to both the base model and the stellar model. As an example, if we take the spectrum of WASP-127 b and we consider the 16 datapoints with wavelengths bluer than 1 m (see Fig. 2), we end up with sixteen absolute distances from the base model and sixteen absolute distances from the stellar model, for a total of thirty-two absolute distances. By considering each spectral Case separately, we then calculate the average of all distances from each model and its uncertainty obtaining an average distance related to the base model () and one referring to the stellar model (). Finally, the stellar activity distance metric (SAD) is defined as the ratio between the two mean absolute distances
@@ Acknowledgements
Funding: This project has received funding from the European Research Council under the European Union Horizon 2020 research and innovation program (grant agreement No. 758892, ExoAI) and the European Union Horizon 2020 COMPET program (grant agreement No. 776403, ExoplANETS A). Furthermore, we acknowledge funding by the UK Space Agency and Science and Technology Funding Council grants ST/K502406/1, ST/P000282/1, ST/P002153/1, ST/ S002634/1, ST/T001836/1, ST/V003380/1, and ST/W00254X/1. This project has received funding from the European Research Council (ERC) under the European Union Horizon 2020 research and innovation programme through Advanced Investigator grant number 883830 (ExoMolHD).
@@
Software: astropy (Astropy Collaboration et al., 2013, 2018; Price-Whelan et al., 2018), Iraclis (Tsiaras et al., 2016b), PyLightcurve (Tsiaras et al., 2016a), ExoTETHyS (Morello et al., 2020), TauREx 3.1 (Al-Refaie et al., 2021), ASteRA (Thompson et al., 2024), YunMa (Ma et al., 2023), h5py (Collette, 2013), emcee (Foreman-Mackey et al., 2013), Matplotlib (Hunter, 2007), Multinest (Feroz et al., 2009), PyMultinest (Buchner et al., 2014), Numpy (Oliphant, 2006), corner (Foreman-Mackey, 2016).